OpenStudy (anonymous):

Select steps that could be used to solve the equation 1 + 3x = -x + 4. a. add x, subtract 1, divide by 4 b. add x, subtract 4, divide by 4 c. subtract 3x, subtract 4, divide by 4 d. subtract 3x, subtract 4, divide by -4 e. subtract 1, add x, divide by 4

OpenStudy (anonymous):

Answer is a :)

OpenStudy (anonymous):

1 + 3x = -x +4 1 + 4x = (x-x) +4 (add x) (1-1) +4x = (4-1) (subtract 1) x= 3/4 (divide by 4 )

OpenStudy (anonymous):

is a the only answer or can anny of the other ones be right to ?

OpenStudy (anonymous):

No :) Thats the only way and can you give medals to who answers ur question :)

OpenStudy (anonymous):

was gettin there and by the way it all of thme excep b,c lol

OpenStudy (anonymous):

Oh :D I didn't looked other answers :D:D:D:D You got me :D

OpenStudy (lilg132):

1 + 3x = -x + 4 you add x which gives you 1 + 3x = 4 and you bring it to the other side 1 + x + 3x = 4 then you minus 1 giving you x + 3x = 4 - 1 which gives you 4x = 3 so then you divide by 4 to leave you with just x 4x / 4 = 3 / 4